Suppose that household incomes in Rock Hill changed from $40,000 to $43,200, and assuming no change in price, the quantity of baloney sold changed from 160 to 150 cases per week. a. What is the value of the income elasticity of demand for baloney? b. Is baloney a normal good or an inferior good?
Answer #1

Answer
Income elasticity of demand=(change in quantity/average quantity)/(change in income/average income)
Change in quantity=150-160=-10
average quantity=(150+160)/2=155
Change in income=43200-40000=3200
average income=(43200+40000)/2=41600
Income elasticity of demand=(-10/155)/(3200/41600)=-0.838709677
=-0.84

the income elasticity of demand for baloney is -0.84

======
b. Is baloney a normal good or an inferior good?
The good is inferior good as the income elasticity of demand is negative. it means an increase in income decreases the demand the good.